Without using tables.find the value of log3.6, if log2=0.3010,log3=0.4771 and log5=0.6990.
Answer by lwsshak3(11628)

Without using tables.find the value of log3.6, if log2=0.3010,log3=0.4771 and log5=0.6990.
..
log 3.6=log 36/10=log36-log10
=log (3^2*2^2)-(log2+log5)
=2log3+2log2-log2-log5
=2log3+log2-log5
=2(.4771)+(.301)-(.699)=.56620
ans:
log 3.6=.56620
Using calculator:
log 3.6=.56630